Three friends are carpooling from city A to city B. Tyrone drives first and goes one fourth of the way. Lisa then takes over, driving two thirds of the remaining distance. John must drive the rest of the way. If the total distance between the cities is 432 miles, then how far does John drive?
How is this calculated? Thank you

Answer by vheroli(126)

Tyrone gets 1/4 of the way so:
432/4=108
432-108=324
There is 324 miles left when Lisa takes over the 2/3 of the way:
324*(2/3)= 216
324-216=108
We are left with 108 miles. Therefore John drive 108 miles
Answer: 108miles